What is the upthrust on a body which displaces
0.6m^3 of water, 0.5m^3 of methylated spirit?
(density of water = 10^3 kgm^-3, density of
methylated spirit = 0.8 x 10^3 kgm^-3 and
g = 10ms^-2)
1
Expert's answer
2019-06-03T09:44:36-0400

Upthrust is an upward force exerted by a fluid that opposes the weight of an immersed body.

The magnitude of the force F is equal to the weight of the fluid that the body displaces. This fluid weight is equal to the density ρ of the fluid multiplied by the volume V and the acceleration due to gravity g


"F=\rho Vg"


First calculate upthrust for water. Substituting "\rho ={{10}^{3}}kg\cdot {{m}^{-}}^{3},\,\, V=0.6{{m}^{3}}" and "g=10m\cdot {{s}^{-}}^{2}" into this formula we get


"F={{10}^{3}}kg\cdot {{m}^{-}}^{3}\cdot 0.6{{m}^{3}}\cdot 10m\cdot {{s}^{-}}^{2}"

"=6\cdot {{10}^{3}}kg\cdot m\cdot {{s}^{-}}^{2}=6\cdot {{10}^{3}}N=6\,kN"

Now calculate upthrust for methylated spirit. Substitute"\rho =0.8\cdot {{10}^{3}}kg\cdot {{m}^{-}}^{3}, \,\,V=0.5{{m}^{3}}"   and "g=10m\cdot {{s}^{-}}^{2}" into formula:


"F=0.8\cdot {{10}^{3}}kg\cdot {{m}^{-}}^{3}\cdot 0.5\,{{m}^{3}}\cdot 10m\cdot {{s}^{-}}^{2}"

"=4\cdot {{10}^{3}}kg\cdot m\cdot {{s}^{-}}^{2}=4\cdot {{10}^{3}}N=4\,kN"

So upthrust for water is "F=6\,kN" , upthrust for methylated spirit is "F=4\,kN"